1. ____ kL = 300 L




PLZ HELP ME!!!
: 1. ____ kL = 300 L




PLZ HELP ME!!!

Answer by stanbon(18994)

____ kL = 300 L
-------------------
The "k" means kilo or one thousand.
300 L = x(1000L)
x = 300/1000
x = 3/10
x = 0.3
--------
Ans: 0.3kL = 300L
=========================
Cheers,
Stan H.